# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ 4. Analysis of I/O Access Patterns
In this section, we aim to present the I/O access patterns commonly found during distributed training. We searched for benchmarks and found DLIO (Devarajan et al., 2021) to be the most promising for simulating the I/O access patterns of DL applications on HPC systems.
@@
The I/O access patterns found during the training phase are influenced by the gradient descent algorithm being used and the training distribution strategy. Here, we describe the I/O access patterns found when using SGD, because SGD is the most popular gradient descent algorithm when training with large datasets. In addition, the workloads presented use data parallelism, where the model is replicated across each worker and the data is partitioned and distributed among them (Alonso-Monsalve et al., 2021). The organization of the samples within files, seen in Fig. 6, has an impact on the I/O operations performed by each worker. Two common organization strategies are one sample per file, commonly used with file formats such as JPEG or WAV, and multiple samples per file, commonly used with file formats such as TFRecord or HDF5. The choice of sample organization plays a large role in the required I/O operations (Lee et al., 2023). A significant amount of additional metadata management overhead can occur when using many files (Aizman et al., 2019). The Lustre PFS used in the workloads presented below has multiple metadata servers (MDS)s which are responsible for providing metadata services such as filename lookup, directory information, file layouts, and access permission. When using one sample per file with a dataset which has a large number of samples, the MDSs can be overwhelmed by concurrent file requests, causing slower training speeds. We show both sample organizations with the Unet3D workload using one sample per file and the BERT workload using multiple samples per file.
@@
We obtained the results below by running DLIO and captured the I/O traces using Darshan (Snyder et al., 2016b). All benchmarks were conducted with Perlmutter’s compute nodes. Perlmutter is a supercomputer located at the National Energy Research Scientific Computing Center of the United States Department of Energy. Each of the compute nodes has physical cores and virtual cores, with a total memory capacity of GiBs. The PFS utilized is an all-flash Lustre file system, with an aggregate bandwidth of over TB/sec and million IOPS ( KiB random). Perlmutter has MDSs and 298 object storage targets (OST)s, and all workloads presented use a MiB stripe size. OSTs are storage devices in which the user’s file data is stored. For visualization, we used dxt-explorer (Bez et al., 2021), which is an interactive web-based tool capable of producing graphs from Darshan DXT logs.
@@
In Fig. 7, we used DLIO to simulate the training of Unet3d, which is used for 3D medical image segmentation tasks. The workload consisted of compute nodes, each running MPI processes. Every MPI process was given access to physical CPU cores, and each TensorFlow process was assigned I/O threads with a batch size of samples. The dataset for Unet3d consists of one sample per NPZ file where each sample is approximately MiBs. There were a total of samples. The model was trained for epochs. Due to the one-to-one correspondence between the number of samples per file, the entire file is requested by the same process at each epoch rather than subsets of the file being requested. In addition, the file was accessed by the same rank ( in this case) at each epoch.
@@
In Fig. 8, we used DLIO to simulate the training of BERT, which is a large language model (LLM). The computational setup consisted of compute nodes, each with MPI processes. Every MPI process was given access to physical CPU cores, and each TensorFlow process was assigned I/O thread. The dataset size was configured to increase training speeds, however, we don’t expect this to significantly change the overall data access patterns. We reduced the number of files from to , and the number of samples per TFRecord file from to . Every sample was bytes. Therefore, there were samples in total. The workload was run twice, once with the original batch size of and a second time with a batch size of . The model was trained for three epochs, which can be clearly seen in the figure.
@@
Fig. 9 is a zoomed-in view of the second epoch of Fig. 8. There are distinct columns of I/O reads that occur at each epoch from every process. This can be expected because the result of would give us the number of batches that need to be read by every process each epoch with multiple samples per file. Taking into account the samples, the samples per batch and the MPI processes, there are batches of samples to read per epoch (rounding up to the nearest integer as the remaining samples need to be read). To ensure this result, we ran the BERT workload again but doubled the batch size to samples. Fig. 10 is a zoomed-in perspective on the second epoch and can be seen to have distinct columns of I/O reads which agrees with batches of samples to read per epoch. The non-overlapping nature of the columns is to be expected due to all processes synchronizing after every batch is read. Since no caching is used for the workloads presented, every read request must reach the underlying PFS for every sample needed. The absence of caching poses a potential I/O bottleneck, particularly when dealing with either large sample sizes or a large number of small samples.
@@ 5.1. Common I/O Optimization Techniques Used by Current ML Frameworks
Here, we explore the data loading and preprocessing capabilities offered by different ML frameworks. We describe the various classes and methods available in state-of-art-frameworks including PyTorch, TensorFlow, Scikit-learn with Dask-ML. Additionally, we touch upon the usage of tools including Horovod for distributed deep learning across various platforms and direct NVIDIA GPU loading for direct loading from storage and network devices to the GPU. The I/O features supported by each ML framework can be seen in Table 4.
@@ 5.1.1. PyTorch
PyTorch uses the Python utility, a popular serialization tool, to save and load the model through the torch.save() and torch.load() methods. PyTorch also has recently (October , in version ) added distributed checkpointing through the torch.distributed.checkpoint API. This allows the loading and saving of models to be done by multiple parallel ranks. It produces at least one file per rank and is capable of checkpointing and loading models in topologies of differing world sizes. To allow the distributed checkpoint to work on differing topologies, the model checkpoint must be resharded. There is no built-in functionality for asynchronous checkpointing; if necessary, the user must implement it.
@@ 5.1.2. TensorFlow
TensorFlow (TensorFlow, 2024) offers the tf.data API to build data loading and preprocessing pipelines. It aims to enable the construction of complex input pipelines from simple reusable pieces. The tf.data.Dataset class is used to represent a sequence of samples. It can be constructed from data stored in memory or from one or more files. If the dataset can fit into node local memory, the Dataset object can be converted to tf.Tensor objects, which can then be fed to the model. Otherwise, the dataset acts as a stream, and data is loaded as needed. The tf.data.Dataset.cache() transformation method can be used to cache the entire dataset in memory or in local storage. However, this method should not be used with large datasets that do not fit in memory or local storage as it will lead to cache thrashing.
@@
TensorFlow has four checkpointing formats: the SavedModel format, HDF5 format, .keras format, and checkpoint format. The SavedModel format stores the entire model graph in a binary file that is then stored within a directory containing a TensorFlow checkpoint. The format is useful for model sharing and deployment with other platforms such as TFLite (TensorFlow, 2024b), TensorFlow.js (Smilkov et al., 2019), and TensorFlowHub (TensorFlow, 2024a). The HDF5 format uses a single HDF5 file and is the legacy choice by TensorFlow when using the higher level tf.keras API. The currently suggested format is the .keras format. It aims to have an efficient name-based saving scheme that enables easier debugging than the legacy (i.e., HDF5) formats. The checkpoint format uses three files: an index file, a metadata file, and a data file. The index file contains an index of the model data, enabling fast retrieval of model values. The metadata file contains the TensorFlow computation graph, which includes the model architecture required to reconstruct the model and the data file contains the model’s actual data (model weights, optimizer states, etc.). It shards the model across multiple files to enable parallelism. Asynchronous checkpointing can be enabled through the tf.train.CheckpointOptions() function. The CheckpointOptions() method takes an enable_async parameter, which, when set to True, moves the checkpointing of the model off the main thread. The return value of tf.train.checkpointOptions() can be passed as a parameter to the tf.train.Checkpoint.restore() and tf.train.Checkpoint.save() methods which are used to load and save the model.
@@ 5.2. Proposed I/O Optimization Techniques
During SGD, samples are commonly shuffled based on a seed using a Pseudo Random Number Generator (PRNG). Each worker is then assigned a partition of the randomly shuffled samples and fetch batches of samples from their assigned dataset partition. The seed can be used to determine the order and frequency at which a given worker will require samples. Dryden et al. (Dryden et al., 2021) propose NoPFS, a machine learning I/O middleware, which capitalizes on these insights to prefetch and cache samples optimally. At the application level, NoPFS can be integrated into the training pipeline of most ML frameworks with minimal code change. NoPFS reduced I/O and improved overall training time by on models using the ImageNet-1k (Deng et al., 2009), ImageNet-22k, and CosmosFlow (Mathuriya et al., 2018) datasets.
Importance sampling is a DNN acceleration method in which samples are skipped based on their “importance” value. The importance value is based on the influence a given sample has on the model’s accuracy. Chen et al. (Chen et al., 2023) introduce iCache, a novel caching approach where samples are placed into one of two in-memory caches: the H-cache or the L-cache. The placement of a sample into either cache is determined by its level of importance, with high-importance samples placed in the H-cache and low-importance samples placed in the L-cache. The technique exploits the fact that when using importance sampling, samples with higher importance values should be fed to the model more frequently than samples with lower importance values. The intelligent caching of samples based on their importance values allows I/O operations to reach the underlying storage system less frequently.
@@
Due to many concurrent processes performing I/O operations during ML training, it can be difficult for middleware to characterize the status of the underlying storage nodes. This makes it challenging to schedule I/O requests to run optimally. Wang et al. (Wan et al., 2020) developed a lightweight ML model to predict the status of the underlying storage nodes and present a case study leveraging their model to effectively balance I/O traffic among the OSSs on a Lustre file system. They tested several models, the most accurate was a support vector machine, which achieved an accuracy of when predicting whether a given OSS was busy or idle.
@@
Various shuffling techniques have been proposed to shuffle samples between workers before each epoch. Global shuffling involves shuffling all samples across all workers, which can make caching and other I/O optimizations difficult. In contrast, local shuffling involves workers being assigned a subset of the dataset, from which they then reuse and reshuffle before each epoch. Hash-based shuffling involves applying a hash function to the keys of the samples, from which the result determines the partition to which the record will be shuffled. Apache Spark uses this technique as a primary shuffling algorithm as it can help achieve load-balancing across partitions and minimize data movement. The randomness of the shuffling does affect the convergence rate of the model (Meng et al., 2019); however, using alternative techniques may allow workers to cache the samples because they no longer read the entire dataset. Partial data shuffling, proposed by Nguyen et al. (Nguyen et al., 2022), achieves similar accuracy to global data shuffling while only requiring each worker to store up to of the dataset. The partial data shuffling algorithm combines the technique of global and local data shuffling. Their paper also discusses that local shuffling achieves accuracy similar to the more common global shuffling strategy in their experimentation. Integrating their solution into PyTorch requires minor modifications of the training script and a load handler of the dataset. Jie et al. (Panayotov et al., 2015b) propose ASRDataset, a system aiming to speed up the training of automatic speech recognition (ASR) models by implementing chunk-level and batch-level shuffling. Their system allows the user to specify how many chunks to batch and the number of samples in each chunk. In addition, ASRDataset uses a high-speed audio data processing library implemented in C++ and CUDA, speeding up audio data augmentation and extraction. Lee and Bahn (Lee and Bahn, 2023) proposed shuffling bundles of samples among workers. This groups samples together to increase locality. To improve buffer caching performance, in each epoch, bundles of samples are accessed in the reverse order of the previous epoch. This improves locality from the buffer cache, allowing the last bundle in the previous epoch to be the first in the next epoch. Liu et al. (Liu et al., 2023b) also propose block shuffling by compacting small files into blocks. They use the ImageNet dataset to test their data-loading pipeline. They cache partially loaded data to avoid redundant reads to the underlying storage system. In the first epoch, all data is cached in memory. All remaining epochs reuse half the loaded data and load the other half from disk to form batches, reducing the underlying storage accesses by half.
@@
Asynchronous I/O enables the overlapping of compute and storage. Sunwoo Lee et al. (Lee et al., 2021) propose an asynchronous I/O strategy to speed up data parallel model training by creating a dedicated I/O thread to run alongside each process. Their model assigns groups of the dataset to training processes and globally shuffles these groups at each epoch. In addition, each process allocates a buffer large enough to store two or more data groups. Each process’s I/O thread reads groups of samples as needed into the buffer, which can occur parallel to model training. The strategy has two drawbacks. The first is the lower degree of randomness due to groups of samples being shuffled, and the second is the increased memory footprint needed to load the additional data groups asynchronously.
@@
Many supercomputing facilities are not yet equipped with large node-local memory devices, which disables the ability to cache datasets in higher storage. Choi et al. (Choi et al., 2023b) aim to enhance I/O performance by creating a DDStore, an in-memory data store. DDStore evenly distributes the chunks in the memory of compute nodes, after which sample accesses from a particular node are made through in-memory read transactions. To reduce possible congestion due to multiple nodes requesting a batch from the same node, DDStore replicates the dataset and distributes it among groups of workers. This allows the samples to be globally shuffled while reducing communication congestion. This strategy makes a large trade-off between reduced node-local memory and fewer PFS I/O operations.
@@
Multiple executing workers can often prepare the same sample redundantly during the training phase. Xie et al. (xie et al., 2022) introduce JOADER, a data loader aimed at reducing the redundant I/O and data preprocessing of samples. JOADER is capable of sharing data preparation work on overlapping datasets. It accomplishes this by dividing the dataset into intersection and difference sets, correlating the selections, and sharing operations that operate on the intersecting set. They also introduce a domain-specific cache policy with a novel tree data structure aimed at effectively caching samples. This strategy is more effective when the preparation of a sample takes a significant amount of time (e.g., image decoding).

### v2

@@ 5. Analysis of I/O Access Patterns
In this section, we aim to present the I/O access patterns commonly found during distributed offline data preparation, training, and inference. We searched for benchmarks and found DLIO (et al, 2021c) to be the most promising to simulate the training I/O access patterns of DL applications in HPC systems.
5.1. Offline Data Preparation Access Patterns
During the offline data preparation phase, I/O access patterns differ depending on the specific operations needed to prepare the dataset for training or inference. Common operations such as data transformations, normalization, or feature scaling can often be distributed across compute nodes. Due to the high complexity that arises in the programming of distributed systems applications, big data frameworks such as Apache Spark (Apache Software Foundation, 2024a), Apache Hadoop (Apache Software Foundation, 2024b), or Dask (Dask, 2024) are often utilized to easily distribute data preparation tasks across nodes. Several popular file formats, including .json, .tfrecord, and .npz, do not natively support random I/O access (Dantas et al., 2022b). Instead, data must be read sequentially, enforcing sequential I/O while eliminating parallelism. These file formats are often ideal for streaming scenarios while hindering dataset preparation operations such as selections and filtering. To enable parallel processing for file formats that do not offer random I/O access, the dataset is often sharded, that is, divided into smaller files. Sharding enables distributed processing among compute nodes on each of the ”shards“ of the dataset. This often results in each process sequentially reading a shard of the dataset, increasing I/O throughput. Although sharding is optional and may require additional processing time, it is often worth the trade-off in reduced I/O times (Aizman et al., 2019). In general, sharding and the use of big data frameworks provide essential tools to enable ideal I/O access patterns and effective use of the HPC system.
5.2. Training I/O Access Patterns
@@
The I/O access patterns found during the training phase are influenced by the gradient descent algorithm used and the training distribution strategy. Here, we describe the I/O access patterns found when using SGD, because SGD is the most popular gradient descent algorithm when training with large datasets. In addition, the presented workloads use data parallelism, where the model is replicated in each worker and the data is partitioned and distributed among them (et al, 2021a). The organization of the samples within the files, as seen in Fig. 7, has an impact on the I/O operations performed by each worker. Two common organizational strategies are one sample per file, often used with file formats such as JPEG or WAV, and multiple samples per file, commonly used with file formats such as TFRecord or HDF5. The choice of sample organization plays a large role in the observed I/O operations (Lee, 2023). A significant amount of additional metadata management overhead can occur when using many files (Aizman et al., 2019). The Luster PFS used in the workloads presented below has multiple metadata servers (MDS)s which are responsible for providing metadata services such as filename lookup, directory information, file layouts, and access permission. When using one sample per file with a dataset with many samples, MDSs can be overwhelmed by concurrent file requests, causing slower training speeds. We show both sample organizations with the Unet3D workload (one sample per file) and the BERT workload (multiple samples per file).
@@
We obtained the results below by running DLIO and captured the I/O traces using Darshan (Snyder et al., 2016). All benchmarks were conducted with Perlmutter’s compute nodes. Perlmutter is a supercomputer located at the National Energy Research Scientific Computing Center of the United States Department of Energy. Each of the compute nodes has physical cores and virtual cores, with a total memory capacity of GiB. The PFS is an all-flash Luster file system, containing a total bandwidth of more than TB/sec and supporting million IOPS ( KiB random). Perlmutter has MDSs and 298 object storage targets (OST)s, and all workloads presented use a MiB stripe size. OSTs are storage devices in which the user’s file data is stored. For visualization, we used DXT Explorer (et al, 2021b), which is an interactive web-based tool capable of producing graphs from Darshan DXT logs.
@@
In Fig. 8, we used DLIO to simulate the training of Unet3D, which is used for 3D medical image segmentation tasks. The workload consisted of compute nodes, each running Python processes. Python processes were given access to physical CPU cores, and the TensorFlow data loader was configured with I/O threads and a batch size of samples. The dataset for Unet3D consists of one sample per NPZ file where each sample is approximately MiB. There were a total of samples. The model was trained for epochs. Due to the one-to-one correspondence between the number of samples per file, the entire file is requested by the same process at each epoch rather than subsets of the file being requested. In addition, the file was accessed by the same rank ( in this case) at each epoch.
@@
In Fig. 9, we used DLIO to simulate the training of BERT, which is a large language model (LLM). The computational setup consisted of compute nodes, each running Python processes. Python processes were given access to physical CPU cores, and the TensorFlow data loader was configured with I/O thread. The dataset size was reduced to increase training speeds; however, we do not expect this to significantly change the overall data access patterns. We reduced the number of files from to , and the number of samples per TFRecord file from to . Each sample was bytes. Therefore, there were samples in total. The workload was run twice, once with the original batch size of and a second time with a batch size of . The model was trained for three epochs.
@@
Fig. 10 is a zoomed-in view of the second epoch of Fig. 9. There are distinct columns of I/O reads that occur at each epoch from every process. This can be expected because the result of would give us the number of batches that need to be read by every process each epoch with multiple samples per file. Taking into account the samples, the samples per batch and the MPI processes, there are batches of samples to read per epoch (rounding up to the nearest integer as the remaining samples need to be read). To ensure this result, we ran the BERT workload again but doubled the batch size to samples. Fig. 11 is a zoomed-in perspective on the second epoch and can be seen to have distinct columns of I/O reads which agrees with batches of samples to read per epoch. The non-overlapping nature of the columns is expected, because all processes synchronize after every batch is read.
Since no caching is used for the workloads presented, every read request must reach the underlying PFS for every sample needed. The absence of caching poses a potential I/O bottleneck, particularly when dealing with large sample sizes or a large number of small samples. The dividing of the datasets among the ranks and the subsequent I/O reads for batches of samples aligns with the observations of other studies documenting ML training I/O access patterns (et al, 2022c, 2021c; Cheng and Gunawi, 2021). The interaction between data organization strategies and the underlying PFS significantly influences the performance, with potential bottlenecks emerging from high volumes of small reads or extensive metadata operations. Understanding and optimizing these patterns is essential for enhancing the efficiency of training ML models on HPC systems.
5.2.1. Checkpoints
As mentioned in Section 3.4, it is common for a consistent pattern of writes to occur during the training loop to checkpoint the model. Both the Unet3D and BERT workloads use the TensorFlow checkpoint function described in Section 6.1.2, where each checkpoint consists of three files: an index file, a metadata file, and a data file. In both workloads, within the checkpoint directory, a file named “checkpoint” has the most recent checkpoint version. For the Unet3D workload, checkpointing was set to occur after epoch , then every epochs thereafter. This resulted in a total of checkpoints after epochs , , and . For BERT, checkpointing was set to occur after every steps. Here, we describe the number of checkpoints when using a batch size of . Since the total number of samples was , the model was trained for epochs, and checkpoints occur every steps, the total number of checkpoints was . As observed, frequent checkpointing can result in a moderate number of files and, consequently, increased write operations. While checkpointing is essential for fault tolerance and recovery, it also introduces additional I/O overhead that may impact training efficiency. ML developers must strike a balance between ensuring node failure resilience through checkpointing and maintaining reasonable I/O performance to avoid training bottlenecks.
5.3. Inference I/O Access Patterns
In this section, we aim to categorize the most common I/O access patterns found during the inference phase within HPC systems. MLPerf (et al, 2020d) developed by MLCommons is a well-known group aiming to provide ML training and inference evaluations. They categorize ML inference into four categories: single-stream, multi-stream, server, and offline (et al, 2020j). However, these inference categories were not specific to HPC environments, and two of the categories (single-stream and server) were not found in the cited literature or in our own experience. We have therefore excluded these categories from this discussion and have found the two categories described below to be representative of most inference configurations found in HPC systems.
Single-stream: multiple client devices stream single sample queries to the HPC system in parallel.
Offline: data is immediately available, and latency and bandwidth are unconstrained, e.g., data is stored in a PFS.
A key difference between streaming and offline inference is the system driving the inference process. In the offline scenario, the HPC system requests data from the storage system, whereas, in streaming scenarios, the HPC system is driven by end users or clients. Latency and bandwidth become more constrained in streaming scenarios, which requires effective load balancing to manage unpredictable query request times to effectively utilize compute nodes (Brewer et al., 2020b, a). A significant amount of research has been done in this area when ML models are deployed on edge computing devices where constraints such as limited memory, high latency, and limited bandwidth are common (Liu et al., 2020; Li et al., 2020; et al, 2022b). Data reduction techniques such as compression (et al., 2020) and intelligent grouping of edge computing resources (Liu et al., 2020) help mitigate network I/O constraints. Although some of the optimizations mentioned above can be used in offline inference, lower latency and higher bandwidth are often found within HPC systems due to high-speed interconnects.
When inference is compared with training, several distinctions emerge. During inference, data is often not sampled randomly, which can result in an I/O access pattern of large contiguous reads in scenarios such as offline inference. High throughput can be expected during offline inference due to PFSs excelling in large contiguous readings. In a streaming scenario, contiguous reads may not be possible if the data is not readily available. Additionally, checkpointing is less common in inference because the model is no longer updated. It should be noted that there are some scenarios, such as online learning and adaptive inference, that may still involve model updates and checkpointing. In addition to checkpointing, storing the prediction results is another potential source of I/O writes. In a streaming scenario, the prediction results may need to be sent back to the querying client, causing more contention on the network bandwidth.
@@ 6.1. Common I/O Optimization Techniques Used by Current ML Frameworks
Here, we explore the data loading and preprocessing capabilities offered by different ML frameworks. We describe the various classes and methods in state-of-the-art frameworks, including PyTorch, TensorFlow, and Scikit-learn with Dask-ML. Additionally, we touch upon the usage of tools, including Horovod for distributed deep learning across various platforms and direct NVIDIA GPU loading for direct loading from storage and network devices to the GPU. The I/O features supported by each ML framework can be seen in Table 4. It should be noted that ML frameworks often aim to be generally applicable and extendable, meaning that it is often possible to extend one framework to have a feature of another. For example, sample caching can be easily integrated into PyTorch even though it does not offer this option through its API. Table 4 only shows a checkmark for a feature if it is easily enabled through the API without extending the underlying framework significantly.
@@ 6.1.1. PyTorch
PyTorch uses the Python utility, a popular serialization tool, to save and load the model through the torch.save() and torch.load() methods. PyTorch also has recently (October , in version ) added distributed checkpointing through the torch.distributed.checkpoint API. This allows the loading and saving of models to be done by multiple parallel ranks. It produces at least one file per rank and is capable of checkpointing and loading models in topologies of differing world sizes. To allow the distributed checkpoint to work on different topologies, the model checkpoint must be resharded. There is no built-in functionality for asynchronous checkpointing; if necessary, the user must implement it.
@@ 6.1.2. TensorFlow
TensorFlow (TensorFlow, 2024) offers the tf.data API (Murray et al., 2021) to build data loading and preprocessing pipelines. It aims to enable the construction of complex input pipelines from simple reusable pieces. The tf.data.Dataset class is used to represent a sequence of samples. It can be constructed from data stored in memory or from one or more files. If the dataset can fit into node local memory, the Dataset object can be converted to tf.Tensor objects, which can then be fed to the model. Otherwise, the dataset acts as a stream, and data is loaded as needed. The tf.data.Dataset.cache() transformation method can be used to cache the entire dataset in memory or in local storage. However, this method should not be used with large datasets that do not fit in memory or local storage as it will lead to cache thrashing.
@@
TensorFlow has four checkpointing formats: the SavedModel format, HDF5 format, .keras format, and checkpoint format. The SavedModel format stores the entire model graph in a binary file that is then stored within a directory containing a TensorFlow checkpoint. The format is useful for model sharing and deployment with other platforms such as TFLite (TensorFlow, 2024b), TensorFlow.js (et al, 2019b), and TensorFlowHub (TensorFlow, 2024a). The HDF5 format uses a single HDF5 file and is the legacy choice by TensorFlow when using the higher level tf.keras API. The currently suggested format is the .keras format. It aims to have an efficient name-based saving scheme that enables easier debugging than the legacy (i.e., HDF5) formats. The checkpoint format uses three files: an index file, a metadata file, and a data file. The index file contains an index of the model data, enabling fast retrieval of model values. The metadata file contains the TensorFlow computation graph, which includes the model architecture required to reconstruct the model and the data file contains the model’s actual data (model weights, optimizer states, etc.). It shards the model across multiple files to enable parallelism. Asynchronous checkpointing can be enabled through tf.train.CheckpointOptions(). This method takes an enable_async parameter, which, when set to True, moves the checkpointing of the model off the main thread. The return value can be passed as a parameter to the tf.train.Checkpoint.restore() and tf.train.Checkpoint.save() methods, which are used to load and save the model.
When performing distributed data parallel training, if the online dataset preparation takes more time than the model training (i.e., time spent performing forward propagation, backpropagation, and sharing of model weights), significant time can be spent waiting for samples to be preprocessed. TensorFlow offers an experimental feature named tf.data.experimental.service (et al, 2023a), which disaggregates the online dataset preparation and training workers. This allows for the ratio of preprocessing workers to training workers to be scaled independently, ensuring that training workers do not have to wait for samples to be preprocessed. The number of online dataset preparation workers and training workers and their CPU/RAM resources can be set either manually or auto-scaled by the service. The service is also capable of sharing preprocessed data between jobs as well as executing coordinated reads to avoid stragglers, which are often caused by the varying input data sizes workers use.
@@ 6.2. Proposed I/O Optimization Techniques
During SGD, samples are commonly shuffled based on a seed using a Pseudo Random Number Generator (PRNG). Each worker is then assigned a partition of the randomly shuffled samples and fetch batches of samples from their assigned dataset partition. The seed can be used to determine the order and frequency at which a given worker will require samples. Dryden et al. (et al, 2021d) propose NoPFS, a machine learning I/O middleware, which capitalizes on these insights to prefetch and cache samples optimally. At the application level, NoPFS can be integrated into the training pipeline of most ML frameworks with minimal code change. NoPFS reduced I/O and improved overall training time by on models using the ImageNet-1k (Deng et al., 2009), ImageNet-22k, and CosmosFlow (et al, 2018b) datasets.
Importance sampling is a DNN acceleration method in which samples are skipped based on their “importance” value. The importance value is based on the influence a given sample has on the model’s accuracy. Chen et al. (et al, 2023b) introduce iCache, a novel caching approach where samples are placed into one of two in-memory caches: the H-cache or the L-cache. The placement of a sample into either cache is determined by its level of importance, with high-importance samples placed in the H-cache and low-importance samples placed in the L-cache. The technique exploits the fact that when using importance sampling, samples with higher importance values should be fed to the model more frequently than samples with lower importance values. The intelligent caching of samples based on their importance values allows I/O operations to reach the underlying storage system less frequently.
@@
Due to many concurrent processes performing I/O operations during ML training, it can be difficult for middleware to characterize the status of the underlying storage nodes. This makes it challenging to schedule I/O requests to run optimally. Wang et al. (et al, 2020k) developed a lightweight ML model to predict the status of the underlying storage nodes and present a case study leveraging their model to effectively balance I/O traffic among the OSSs on a Lustre file system. They tested several models, the most accurate was a support vector machine, which achieved an accuracy of when predicting whether a given OSS was busy or idle.
@@
Various shuffling techniques have been proposed to shuffle samples between workers before each epoch. Global shuffling involves shuffling all samples across all workers, which can make caching and other I/O optimizations difficult. In contrast, local shuffling involves workers being assigned a subset of the dataset, from which they then reuse and reshuffle before each epoch. Hash-based shuffling involves applying a hash function to the keys of the samples, from which the result determines the partition to which the record will be shuffled. Apache Spark uses this technique as a primary shuffling algorithm as it can help achieve load-balancing across partitions and minimize data movement. The randomness of the shuffling does affect the convergence rate of the model (Meng et al., 2019); however, using alternative techniques may allow workers to cache the samples because they no longer read the entire dataset. Partial data shuffling, proposed by Nguyen et al. (et al, 2022c), achieves similar accuracy to global data shuffling while only requiring each worker to store up to of the dataset. The partial data shuffling algorithm combines the technique of global and local data shuffling. Their paper also discusses that local shuffling achieves accuracy similar to the more common global shuffling strategy in their experimentation. Integrating their solution into PyTorch requires minor modifications of the training script and a load handler of the dataset. Jie et al. (Jie et al., 2023) propose ASRDataset, a system aiming to speed up the training of automatic speech recognition (ASR) models by implementing chunk-level and batch-level shuffling. Their system allows the user to specify how many chunks to batch and the number of samples in each chunk. In addition, ASRDataset uses a high-speed audio data processing library implemented in C++ and CUDA, speeding up audio data augmentation and extraction. Lee and Bahn (Lee and Bahn, 2023) proposed shuffling bundles of samples among workers. This groups samples together to increase locality. To improve buffer caching performance, in each epoch, bundles of samples are accessed in the reverse order of the previous epoch. This improves locality from the buffer cache, allowing the last bundle in the previous epoch to be the first in the next epoch. Liu et al. (Liu et al., 2023) also propose block shuffling by compacting small files into blocks. They use the ImageNet dataset to test their data-loading pipeline. They cache partially loaded data to avoid redundant reads to the underlying storage system. In the first epoch, all data is cached in memory. All remaining epochs reuse half the loaded data and load the other half from disk to form batches, reducing the underlying storage accesses by half.
@@
Asynchronous I/O enables the overlapping of compute and storage. Sunwoo Lee et al. (et al, 2021f) propose an asynchronous I/O strategy to speed up data parallel model training by creating a dedicated I/O thread to run alongside each process. Their model assigns groups of the dataset to training processes and globally shuffles these groups at each epoch. In addition, each process allocates a buffer large enough to store two or more data groups. Each process’s I/O thread reads groups of samples as needed into the buffer, which can occur parallel to model training. The strategy has two drawbacks. The first is the lower degree of randomness due to groups of samples being shuffled, and the second is the increased memory footprint needed to load the additional data groups asynchronously. Serizawa et al. (Serizawa et al., 2019) also propose overlapping the copying and reading of the dataset for training. They observed to times speed up when using their asynchronous strategy compared to using Lustre directly.
@@
Many supercomputing facilities are not yet equipped with large node-local memory devices, which disables the ability to cache datasets in higher storage. Choi et al. (et al, 2023e) aim to enhance I/O performance by creating a DDStore, an in-memory data store. DDStore evenly distributes the chunks in the memory of compute nodes, after which sample accesses from a particular node are made through in-memory read transactions. To reduce possible congestion due to multiple nodes requesting a batch from the same node, DDStore replicates the dataset and distributes it among groups of workers. This allows the samples to be globally shuffled while reducing communication congestion. This strategy makes a large trade-off between reduced node-local memory and fewer PFS I/O operations. Dantas et al. (Dantas et al., 2022a) created Monarch, which aims to exploit storage tiering to enable fast training speeds on single-node workloads. Their solution aims to be ML framework agnostic acting as a transparent middleware which intelligently caches at different storage layers that may not fit into host memory. Aizman et al. (Aizman et al., 2019) created a storage system named AIStore aiming to provide an infinitely scalable namespace. The namespace can be used on an arbitrary number of disks. Metadata overhead is minimized by enabling dataflow directly between compute nodes and storage targets.
@@
Multiple executing workers can often prepare the same sample redundantly during the training phase. Xie et al. (et al, 2022f) introduce JOADER, a data loader aimed at reducing the redundant I/O and data preprocessing of samples. JOADER is capable of sharing data preparation work on overlapping datasets. It accomplishes this by dividing the dataset into intersection and difference sets, correlating the selections, and sharing operations that operate on the intersecting set. They also introduce a domain-specific cache policy with a novel tree data structure aimed at effectively caching samples. This strategy is more effective when the preparation of a sample takes a significant amount of time (e.g., image decoding). Jin et al. (Jin et al., 2024) created MMDataLoader, a data preprocessing pipeline framework aiming to increase preprocessing throughput and reduce redundant data preparation. It operates at the server level, allowing it to share results across training tasks. The framework enables batches of data to be preprocessed for multiple models at the same time, eliminating the need for each task to recompute batches individually.